# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ VI-C Evaluation Metrics
1) Evaluation Metrics for Node Level Tasks: Different from static graph signal processing, in the time-varying graph, the most popular task is node-level regression rather than classification, as the core in TV-GSP is the modeling to graph temporal association. Thus, learning from the past to make accurate predictions of the future is the property that most TV-GSP models desire. The metrics for regression are usually drawn from multi-variate time series analysis, which mainly include Mean-Square Error (MSE), Mean Absolute Error (MAE) [118], root Normalized Mean-Square Error(rNMSE) [179], Root Mean Squared Error (RMSE) [180] and Mean Absolute Percentage Error (MAPE) [181]. Taking step ahead prediction as an example, these metrics are defined as functions of node signal as:
@@
In the offline GSP algorithms, they also get to work by replacing the L2 norm with the Frobenius norm, and the functions are then built between node signal matrix and estimated node signal matrix . In addition to regression, there are also some node-level classification tasks on time-varying graphs that usually use Accuracy, Precision, Recall, Area Under Curve for binary classification, and Micro-F1 and macro-f1 for multi-class classification [122, 118, 182, 183]. The formulation of these metrics is the same as seen in static graph signal classification and conventional machine learning tasks.
@@
2) Evaluation Metrics for Edge and Graph Level Tasks: Besides node-level tasks, there are also some edge-level and graph-level tasks on time-varying graphs (e.g., edge prediction and graph classification). Note that, in an unweighted graph, the edge prediction can be viewed as edge classification (i.e., to judge whether the edge exists in the future or not). Thus, the metrics that are popular in classification can also be utilized for them [184, 155, 185]. Besides, for graph topology regression, the regression error of estimated graph topology can be computed by comparing the graph adjacency matrix. In this situation, the input to the regression metrics functions defined in Eq. 31 becomes the adjacency matrix, and the L2 norm becomes Fronbenius norm [186].
@@
3) Evaluation Metrics for Graph Lifelong Learning: In graph lifelong learning, the evaluation metrics change slightly, as its goal is to maintain the ability to make predictions on previous nodes/classes. In this way, the metrics that are usually used include “AP” and “AF” which represent average performance (e.g., the average classification accuracy or regression error) and average forgetting rate. They can be computed by:
@@
VII Conclusion and Future Directions
Time-varying graph signal processing represents a significant advancement in the analysis of dynamic systems. By extending traditional GSP techniques to account for temporal variations in graph structure and node signals, TV-GSP provides a powerful framework for analyzing a wide range of real-world systems. As the field continues to evolve, the integration of TV-GSP with topological signal processing, spatiotemporal graph neural networks, and other advanced techniques will open up new possibilities for research and application.
The field of TV-GSP is still in its infant stage, with many open challenges and opportunities for future research. Some key directions include:
Scalability: As the size of graphs and the length of time series data increase, scalable algorithms that can handle large-scale time-varying graphs are needed. This requires developing efficient computational methods that reduce the complexity of TV-GSP without sacrificing accuracy.
Real-Time Processing: Many applications require real-time analysis of time-varying graphs, such as in cybersecurity or financial markets. Future research should focus on developing real-time TV-GSP methods that can adapt quickly to changes in the graph structure and node signals.
Integration with Machine Learning: There is significant potential in integrating TV-GSP with machine learning techniques, particularly in the context of deep learning. Developing hybrid models that combine the interpretability of GSP with the predictive power of machine learning could lead to new insights and applications.
Theoretical Foundations: Theoretical advancements are needed to better understand the behavior of TV-GSP techniques, especially in terms of stability, convergence, and robustness. Establishing a solid theoretical framework will help guide the development of new methods and applications.
Applications: Finally, exploring new applications of TV-GSP in areas such as neuroscience, epidemiology, and smart cities will drive the field forward. These applications can benefit from the ability to analyze complex, time-varying data on graphs. Conclusion

### v2

@@ Appendix B Evaluation Metrics
Different from static GSP, in TVGs, the most popular task is node-level regression rather than classification, as the core in TVGSP is the modeling of graph temporal association. Thus, learning from the past to make accurate predictions of the future is the property that most TVGSP models desire. The metrics for regression are usually drawn from multivariate time series analysis, which mainly include Mean-Square Error (MSE), Mean Absolute Error (MAE) [180], root Normalized Mean-Square Error(rNMSE) [153], Root Mean Squared Error (RMSE) [3], and Mean Absolute Percentage Error (MAPE) [152]. Taking step ahead prediction as an example, these metrics are defined as functions of node signal as:
@@
In GSP and graph machine learning algorithms, the Frobenius norm serves as an alternative to the norm, where functions are constructed between the node signal matrix and the estimated node signal matrix . In addition to regression, there are also classification tasks on TVGs that usually use Accuracy, Precision, Recall, Area Under Curve for binary classification, and micro-F1 and macro-F1 for multi-class classification [90, 180, 147, 193]. The formulation of these metrics is the same as in static graph signal classification and conventional machine learning tasks. Table 2 illustrates the performance of various TVGSP models on a 1-step ahead spatio-temporal prediction task. The methods being compared are from 2 categories: online GSP methods GLMS [31], GNLMS [170], and GNS [134], as well as offline TVGNN models such as STGCN [207], and RGDAN [41]. We included the GCN [86] as the non-temporal GNN baseline. These models are evaluated for wind speed and temperature prediction using the RMSE and MAE.
@@
Besides node-level tasks, there are also some edge-level and graph-level tasks on TVGs (e.g., edge prediction and graph classification). Note that, in an unweighted graph, the edge prediction can be viewed as edge classification (i.e., to judge whether the edge exists in the future or not). Thus, the metrics that are popular in classification can also be utilized for them [214, 215, 44]. Besides, for graph topology regression, the regression error of the estimated graph topology can be computed by comparing the graph adjacency matrix. In this situation, the input to the regression metrics functions defined in (54) becomes the adjacency matrix, and the L2 norm becomes the Frobenius norm [135].
@@
In graph lifelong learning, the evaluation metrics change slightly, as its goal is to maintain the ability to make predictions on previous nodes/classes. In this way, the metrics that are usually used include "AP" and "AF," which represent average performance (e.g., the average classification accuracy or regression error) and average forgetting rate. They can be computed by:
@@
| Notation | Meaning |
| Graph | |
| Node set | |
| Edge set | |
| Number of nodes | |
| Graph signal at time | |
| Adjacency matrix | |
| Laplacian matrix | |
| Eigenvector matrix of | |
| Eigenvalue matrix of | |
| Spectral filter | |
| Weights | |
| Sampling / masking matrix | |
| Loss function | |
| Temporal Change | |
| TVG | Time-Varying Graph |
| STG | Spatiotemporal Graph |
| DTDG | Discrete-Time Dynamic Graph |
| CTDG | Continuous-Time Dynamic Graph |
| STGNN | Spatiotemporal Graph Neural Network |
| DGNN | Dynamic Graph Neural Network |
| GFT | Graph Fourier Transform |
| JFT | Joint Time–Graph Fourier Transform |